Tues 6/9
3 yrly electrical survey completed.
Bilge float switch port side replaced and wire damage repaired.
Stbd a/c control unit replaced.
Wed 8/9
Sunday run at 1400 gl4 - GL2 peggy's 4/5 indefinitely. I have entered a reminder into S planner and note on helm.
0540 run GL2-GL4 cancelled immediately check emails
Thur 9/9
200hr genset svc
14/9 Both mains 250hr svc ...check anodes, all ok.
New page section 4 sms... ukc 0.75 on chart datum.
Check Dave's email....backup for Torresian 0616 and 1700, wed, thur and fri. Expecting extra pax re shutdown crew. This will necessitate one crew to start early or finish late. Work it out and call Dave.
Genset service overdue....possibly booked in thru Dave for Wed 1300 or later....check engineering.
Any questions give me a call.